OpenStudy (anonymous):

2. An electron travels at a speed of 2.0 x 10^4 m/s through a uniform magnetic field whose magnitude is 1.2 x 10^-3 T. What is the magnitude of the magnetic force on the electron if its velocity vector and the magnetic field vector (a) are perpendicular, (b) make an angle of 45 degrees, or (c) are parallel?

OpenStudy (anonymous):

f=B*q*v*sin @. here B=magnitude of the magnetic field. q=charge of the electron=1.6*10^-19 coulombs v=velocity of the electron. @=angle btw magnetic field and the velocity of the electron.